OpenStudy (anonymous):

can someone explain the steps to this problem.... Differentiate the the function F(x)=2 ^40

OpenStudy (anonymous):

Is there no independent variable (x-value, maybe) in there?

OpenStudy (anonymous):

no my book has it as f(x)=2^40 but the answer to this is zero i have no idea how they got that just need the steps explained

OpenStudy (anonymous):

constants go to zero automatically

sam (.sam.):

F(x)=2^(40) The derivative of F(x) is equal to F'(x)=(d)/(dx) 2^(40). F'(x)=(d)/(dx) 2^(40) The derivative of a constant with respect to x is 0. 0 The derivative of the function is 0. F'(x)=0

OpenStudy (anonymous):

Okay, I'll ask you this question. When you have y = 2, what do you have? A horizontal line at 2. What is the slope? Zero. If you have y = a number, you'll always get that horizontal line and your slope will always be zero.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome